# Federated Multi-armed Bandits Under Byzantine Attacks

### v2

@@ III-C Regret
IV Fed-MoM-UCB Algorithm
We propose Fed-MoM-UCB: Federated Median-of-Means Upper Confidence Bound algorithm to minimize the robust federated regret under Byzantine clients. In Fed-MoM-UCB, the clients roughly follow the same arm sampling rule as in Fed2-UCB in [21]. However, the server employs an entirely different MoM-based robust estimator for global arm outcomes, which constitutes the key component of Fed-MoM-UCB. The server communicates with the clients at the end of every phase. Each phase lasts for a certain number of rounds communicated to the clients by the server. Next, we explain in detail client and server sides of our distributed algorithm.
IV-A Client Side
@@ IV-B Server Side
Let $\alpha:[0,1/2]\rightarrow[0,1]$ be a mapping such that $\forall\lambda\in[0,1/2),~{}2\lambda<\alpha(\lambda)<1$. Table II lists examples of $\alpha$.
TABLE II: Different choices $\alpha(\lambda)$ satisfying Definition 1.
@@
Since our primary objective is to identify a set of $\beta$-suboptimal arms, our analysis concentrates on the algorithm’s regret after it has completed the elimination of $\beta$-suboptimal arms and all communication rounds have concluded. We refer to this state as the termination phase of our algorithm. In Section V-B, we show that Fed-MoM-UCB successfully reaches the termination phase, and provide sample complexity and regret bounds at this point. More specifically, we denote the termination phase of our algorithm as $p_{\tau}$ and denote the regret at termination as
@@ V Theoretical Analysis
V-A Preliminary Technical Results
@@
Proof.
The proof builds upon the proof of MoM concentration bound given in [48, Proposition 2]. Let ${\mathcal{G}}_{p}$ denote the set of client groups formed by the server at the end of phase $p$ and ${\mathcal{G}}^{\text{tru}}_{p}\subseteq{\mathcal{G}}_{p}$ denote the subset of groups with no Byzantine clients. We denote by $G_{p}:=\lvert{\mathcal{G}}_{p}\rvert$ and $G^{\text{tru}}_{p}:=|{\mathcal{G}}^{\text{tru}}_{p}|$ the cardinalities of these sets. There is no concentration guarantee for the groups that include Byzantine clients. The worst case is when each Byzantine client belongs to a different group, rendering the maximum number of groups unreliable. Since $M^{\text{mal}}\leq\lambda M$ due to Assumption 1, at most $\lfloor\lambda M\rfloor$ of the $G_{p}$ groups will be corrupted by Byzantine clients. Then, we have,
$\mathbb{P}(|\overline{U}_{k}(p)-\mu_{k}|>\gamma_{p}|{\mathcal{G}}_{p})$
By definition, Term I is the average of $B$ independent $\sigma/\sqrt{s(p)}$-subgaussian random variables. Thus, Term I is $\sigma/\sqrt{s(p)B}$-subgaussian. Similarly, Term II is the average of $B$ independent $\sigma_{c}$-subgaussian random variables. Therefore, it is $\sigma_{c}/\sqrt{B}$-subgaussian. As a result, Term I plus Term II is $\sigma/\sqrt{s(p)B}+\sigma_{c}/\sqrt{B}$-subgaussian. Let $\rho\coloneqq\sigma/\sqrt{s(p)}+\sigma_{c}$ and $\rho^{\prime}\coloneqq\frac{1}{\sqrt{B}}\rho$. Then, using the subgaussian tail inequality in [49, Theorem 5.3 ]), we have
where (14) follows from (8), and (15) follows since $\eta(\lambda)>1/2$ implies that $6\eta(\lambda)-1>2\eta(\lambda)$. From (12), observe that (15) holds and $\eta(\lambda)G_{p}-G_{p}^{\text{tru}}\leq 0$, which gives
(12)$\leq q_{p}^{\frac{2\eta(\lambda)-1}{2}G_{p}}2^{\eta(\lambda)G_{p}}~{}.$ (16)
for all $\gamma_{p}^{2}\geq\frac{4\rho^{2}}{B}\frac{4\eta(\lambda)-1}{2\eta(\lambda)-1}\log 2$, which completes the proof. ∎
Lemma 1 hints on how to choose $E_{p}$ in (6) and $M$. However, observe that it is not possible to provide any guarantees for an arbitrarily small uncertainty level $\gamma_{p}$ in (9). The smallest uncertainty level under which (9) holds is equal to the square root of the term that appears on the r.h.s. of (8). Therefore, we set the uncertainty level $E_{p}$ as
@@
For all $\alpha(\lambda)$ in Table II except arithmetic mapping, $\lim_{\lambda\rightarrow 0}\omega=0$.
@@
Next, we derive how many clients to recruit such that when $E_{p}$ is given as in (17), (6) holds. Let
@@
Proof.
where (22) follows from the definitions of $\rho$ and $E_{p}$. Finally, observe that our choice of $M$ satisfies (22). ∎
In the rest of the paper, we always assume that $E_{p}$ is set as in (17) and $M$ as in Lemma 3.
@@
V-B $(\beta,\delta)$-PACness, Sample Complexity and Regret Analysis
@@
The following lemma shows that all $\beta$-suboptimal arms are eliminated by termination phase.
@@ VI Experimental Results
where arms are $\varrho$-subgaussian. For our setting we pick $\varrho=\sigma+\sigma_{c}$. Similar to Fed-MoM-UCB, we set the termination condition for Fed-Trim-UCB to $E_{p,\text{trim}}\leq\beta/4$ and use the same client recruiting policy. For our analysis, we conducted simulations on a synthetic MAB problem, applying various attack schemes to evaluate the efficacy of Fed-MoM-UCB, Fed-Trim-UCB, and Fed2-UCB. We detail our findings across three distinct scenarios, with each scenario’s objectives and results thoroughly discussed in its designated section. For a concise overview of these scenarios, see Table IV The model and arm sampling noises are defined as zero-mean additive i.i.d. Gaussian noise with variance $\sigma_{c}^{2}=0.02^{2}$ and $\sigma^{2}=2^{2}$ for all scenarios. We use $\delta=0.001$ and perform 100 independent experiments and report the mean and standard deviations for the cumulative regret and the communication cost. We calculate the cumulative regret using (3) for all of the algorithms, where the $\beta$-optimal arms do not incur any regret.
TABLE III: Unknown global mean arm outcomes
@@
TABLE IV: Parameters for different Scenarios. $M$, $B$, and $G$ columns show the number of total clients recruited, the number of clients in a group, and the number of groups for the Fed-MoM-UCB algorithm, respectively.
@@
VI-A Scenario 1
We choose $\beta=0.32$ which gives the $\beta$-suboptimal arm set as $\neg{\cal K}_{\beta}=\{1,\ldots,33\}$. For this scenario we consider a case where the fraction of Byzantine clients is small but the attacking scheme is strong. We set the Byzantine clients’ fraction to $\lambda=0.05$ and use the geometric mapping for Fed-MoM-UCB, that is, $\alpha(\lambda)=\sqrt{2\lambda}$ (see Table II). We choose the geometric mapping as it yields a small value for the discernibility margin at $\lambda=0.05$, which is $\omega=0.149$ while also resulting in small constants $\Lambda_{1}$ and $\Lambda_{2}$ that scale the number of clients to be recruited and the cumulative regret. The communication cost is set to $C=1$. Finally, we consider the following outlier attack scheme for a Byzantine client $m$,
@@ VI-A Scenario 1
Unlike Fed-MoM-UCB, Fed2-UCB continues to recruit more clients over phases to reduce the uncertainty emerging from the client sampling. As discussed in Section V, that does not help in the presence of Byzantine clients. We set $s(p)=10p$ and run Fed2-UCB with two different configurations, $M(p)=10p$ and $M(p)=2^{p}$, where $M(p)$ is the total number of clients recruited at phase $p$. Fig. 2(d) confirms that Fed2-UCB terminates. That is, only a single active arm is left, and the communication is stopped. However, the remaining arm is not a $\beta$-optimal arm and Fed2-UCB incurs linear regret in both cases as can be seen in Fig. 2(a) and Fig. 2(b). Moreover, while recruiting more clients does not help Fed2-UCB eliminate the $\beta$-suboptimal arms, the cumulative regret also explodes since the more clients are recruited, the more the suboptimal arms are played. Table IV shows the extreme difference between the total number of clients recruited by Fed2-UCB and by Fed-MoM-UCB when they terminate. The unlimited client recruitment scheme of Fed2-UCB does not only lead the cumulative regret to explode in the presence of Byzantine clients, but it may also not be feasible for some applications. Fed-MoM-UCB recruits $M=126$ only once at the beginning and creates $G=42$ groups. Note that even though it is possible to have $B\geq 4$ as well while still maintaining the robustness, the geometric mapping is the best option for $\lambda=0.05$ since it gives the small $\omega$, $\Lambda_{1}$, and $\Lambda_{2}$ terms. We run Fed-MoM-UCB with two different configurations, $s(p)=10p$ and $s(p)=2^{p}$ as in Corollary 1. Fig. 2(d) confirms that Fed-MoM-UCB terminates and stops the communication in both cases. From Fig. 2(b), we can see that Fed-MoM-UCB is able to eliminate all the $\beta$-suboptimal arms as the cumulative regret cease to increase. We run Fed-Trim-UCB with $s(p)=10p$. Similar to Fed-MoM-UCB, Fed-Trim-UCB also achieves to eliminate all the $\beta$-suboptimal arms. However, as can be seen from Fig. 2 the elimination requires more samples compared to Fed-MoM-UCB, hence the cumulative regret is greater. Also note that while Fed-Trim-UCB manages to eliminate all $\beta$-suboptimal arms, it fails to terminate in $5\times 10^{4}$ rounds, hence the communication cost is higher. This is because the confidence bound for Fed-Trim-UCB decreases slower compared to Fed-MoM-UCB.
VI-B Scenario 2
@@
VI-C Scenario 3
[…]
Next, we analyze the mappings $\alpha(\lambda)$ given in Table II. Calculating $\omega$ for each choice, we observe that the smallest $\omega$ value is obtained with the geometric mapping which also accumulates the smallest regret and terminates the quickest. While due to its small number of client recruitment, the arithmetic mapping incurs small regret, it results in the largest number of rounds necessary to terminate. This is because while the small group size resulting from the arithmetic mapping gives rise to a small $M$, it also increases the number of samples required for elimination of $\beta$-suboptimal arms. Fig. 4(a) shows, the order between the cumulative regrets of the mappings as well as the communication costs. The geometric mapping offers the best performance with $M=92$ client, while the polynomial mapping offers the worst with the maximum number of clients $M=288$. Fig. 4(b) shows all mappings result in the termination of the algorithm with the geometric mapping giving the quickest termination and the arithmetic mapping giving the slowest.
VII Conclusion

### v3

@@ 3.3 Regret
4 Fed-MoM-UCB Algorithm
We propose Fed-MoM-UCB: Federated Median-of-Means Upper Confidence Bound algorithm for $\beta$-optimal arm identification under Byzantine clients. In addition, Fed-MoM-UCB will also guarantee small robust federated regret. It is a distributed algorithm that consists two pieces: one piece running independently on each client, and one piece running on the server. In Fed-MoM-UCB, the clients roughly follow the same arm sampling rule as in Fed2-UCB in [24]. However, the server employs an entirely different MoM-based robust estimator for global arm outcomes, which constitutes the key component of Fed-MoM-UCB. The server communicates with the clients at the end of every phase. Each phase lasts for a certain number of rounds communicated to the clients by the server. Next, we explain in detail client and server sides of our distributed algorithm.
4.1 Client Side
@@ 4.2 Server Side
Let $\alpha:[0,1/2]\rightarrow[0,1]$ be a mapping such that $\forall\lambda\in[0,1/2),~{}2\lambda<\alpha(\lambda)<1$. Table 2 lists examples of $\alpha$.
Table 2: Different choices $\alpha(\lambda)$ satisfying Definition 1.
@@
Since our primary objective is to identify a set of $\beta$-optimal arms, our analysis concentrates on the algorithm’s regret after it has completed the elimination of $\beta$-suboptimal arms and all communication rounds have concluded. We refer to this state as the termination phase of our algorithm. In Section 5.2, we show that Fed-MoM-UCB successfully reaches the termination phase, and provide sample complexity and regret bounds at this point. More specifically, we denote the termination phase of our algorithm as $p_{\tau}$ and denote the regret at termination as
@@ 5 Theoretical Analysis
5.1 Preliminary Technical Results
@@
The proof of Lemma 1 utilizes the fact that a mapping $\alpha(\lambda)$ as defined in Definition 1, ensures that at least half of the groups consists solely of honest clients. The complete proof in detail is provided in the Appendix. Lemma 1 offers guidance on selecting $E_{p}$ in (6) and $M$, while also highlighting a limitation: it is not possible to achieve probability guarantees for an arbitrarily small uncertainty level $\gamma_{p}$ in (9). The smallest uncertainty level under which (9) holds is equal to the square root of the term that appears on the r.h.s. of (8). Therefore, we set the uncertainty level $E_{p}$ as
@@
For all $\alpha(\lambda)$ in Table 2 except arithmetic mapping, $\lim_{\lambda\rightarrow 0}\omega=0$.
@@
Lemma 2 is based on the fact that as the sampling budget $s(p)$ increases, the uncertainty level $E_{p}$ decreases and approaches $\omega/4$ as $p$ grows. Consequently, for $\beta>\omega$, there exists a phase $p$ at which the termination condition $E_{p}\leq\beta/4$ in Algorithm 1 is met. Next, we derive how many clients to recruit such that when $E_{p}$ is given as in (10), (6) holds. Let
@@
The proof of Lemma 3, provided in the Appendix, combines the confidence bound (9) from Lemma 1 with a union bound over all phases $p\in[\bar{p}]$. Then, the number of clients $M$ is chosen as the smallest value that ensures $P({\cal E}_{g})\geq 1-\delta$. Choosing a larger $M$ is unnecessary, as it does not reduce the discernibility margin, as discussed in Remark 2. Additionally, the fixed client selection scheme used in Fed-MoM-UCB is simpler to implement than schemes that increase the number of clients with each phase, as in [24]. In the rest of the paper, we always assume that $E_{p}$ is set as in (10) and $M$ as in Lemma 3.
@@
5.2 $(\beta,\delta)$-PACness, Sample Complexity and Regret Analysis
@@
Lemma 6 implies that arms that have a suboptimality gap $\Delta_{k}$ which is closer to the discernibility margin requires more samples to eliminate. The following lemma shows that all $\beta$-suboptimal arms are eliminated by termination phase.
@@ 6 Experimental Results
where arms are $\varrho$-subgaussian. For our setting we pick $\varrho=\sigma+\sigma_{c}$. Similar to Fed-MoM-UCB, we set the termination condition for Fed-Trim-UCB to $E_{p,\text{trim}}\leq\beta/4$ and use the same client recruiting policy. For our analysis, we conducted simulations on a synthetic MAB problem, applying various attack schemes to evaluate the efficacy of Fed-MoM-UCB, Fed-Trim-UCB, and Fed2-UCB. We detail our findings across three distinct scenarios, with each scenario’s objectives and results thoroughly discussed in its designated section. For a concise overview of these scenarios, see Table 4. The model and arm sampling noises are defined as zero-mean additive i.i.d. Gaussian noise with variance $\sigma_{c}^{2}=0.02^{2}$ and $\sigma^{2}=2^{2}$ for all scenarios. We use $\delta=0.001$ and perform 100 independent experiments and report the mean and standard deviations for the cumulative regret and the communication cost. We calculate the cumulative regret using (3) for all of the algorithms, where the $\beta$-optimal arms do not incur any regret.
Table 3: Unknown global mean arm outcomes
@@
Table 4: Parameters for different Scenarios. $M$, $B$, and $G$ columns show the number of total clients recruited, the number of clients in a group, and the number of groups for the Fed-MoM-UCB algorithm, respectively.
@@
6.1 Scenario 1
We choose $\beta=0.32$ which gives the $\beta$-suboptimal arm set as $\neg{\cal K}_{\beta}=\{1,\ldots,33\}$. For this scenario we consider a case where the fraction of Byzantine clients is small but the attacking scheme is strong. We set the Byzantine clients’ fraction to $\lambda=0.05$ and use the geometric mapping for Fed-MoM-UCB, that is, $\alpha(\lambda)=\sqrt{2\lambda}$ (see Table 2). We choose the geometric mapping as it yields a small value for the discernibility margin at $\lambda=0.05$, which is $\omega=0.149$ while also resulting in small constants $\Lambda_{1}$ and $\Lambda_{2}$ that scale the number of clients to be recruited and the cumulative regret. The communication cost is set to $C=1$. Finally, we consider the following outlier attack scheme for a Byzantine client $m$,
@@ 6.1 Scenario 1
Unlike Fed-MoM-UCB, Fed2-UCB continues to recruit more clients over phases to reduce the uncertainty emerging from the client sampling. As discussed in Section 5, that does not help in the presence of Byzantine clients. We set $s(p)=10p$ and run Fed2-UCB with two different configurations, $M(p)=10p$ and $M(p)=2^{p}$, where $M(p)$ is the total number of clients recruited at phase $p$. Fig. 3(d) confirms that Fed2-UCB terminates. That is, only a single active arm is left, and the communication is stopped. However, the remaining arm is not a $\beta$-optimal arm and Fed2-UCB incurs linear regret in both cases as can be seen in Fig. 3(a) and Fig. 3(b). Moreover, while recruiting more clients does not help Fed2-UCB eliminate the $\beta$-suboptimal arms, the cumulative regret also explodes since the more clients are recruited, the more the suboptimal arms are played. Table 4 shows the extreme difference between the total number of clients recruited by Fed2-UCB and by Fed-MoM-UCB when they terminate. The unlimited client recruitment scheme of Fed2-UCB not only causes cumulative regret to increase significantly in the presence of Byzantine clients but may also be impractical for certain applications. In contrast, Fed-MoM-UCB achieves the $\beta$-optimal arm identification goal with a much smaller, fixed number of clients. In real-world applications, such as the treatment evaluation example in Section 2.3, progressively recruiting additional clients may be unrealistic.
Fed-MoM-UCB recruits $M=126$ only once at the beginning and creates $G=42$ groups. Note that even though it is possible to have $B\geq 4$ as well while still maintaining the robustness, the geometric mapping is the best option for $\lambda=0.05$ since it gives the small $\omega$, $\Lambda_{1}$, and $\Lambda_{2}$ terms. We run Fed-MoM-UCB with two different configurations, $s(p)=10p$ and $s(p)=2^{p}$ as in Corollary 1. Fig. 3(d) confirms that Fed-MoM-UCB terminates and stops the communication in both cases. From Fig. 3(b), we can see that Fed-MoM-UCB is able to eliminate all the $\beta$-suboptimal arms as the cumulative regret cease to increase. We run Fed-Trim-UCB with $s(p)=10p$. Similar to Fed-MoM-UCB, Fed-Trim-UCB also achieves to eliminate all the $\beta$-suboptimal arms. However, as can be seen from Fig. 3 the elimination requires more samples compared to Fed-MoM-UCB, hence the cumulative regret is greater. Also note that while Fed-Trim-UCB manages to eliminate all $\beta$-suboptimal arms, it fails to terminate in $5\times 10^{4}$ rounds, hence the communication cost is higher. This is because the confidence bound for Fed-Trim-UCB decreases slower compared to Fed-MoM-UCB.
6.2 Scenario 2
@@
6.3 Scenario 3
[…]
Next, we analyze the mappings $\alpha(\lambda)$ given in Table 2. Calculating $\omega$ for each choice, we observe that the smallest $\omega$ value is obtained with the geometric mapping which also accumulates the smallest regret and terminates the quickest. While due to its small number of client recruitment, the arithmetic mapping incurs small regret, it results in the largest number of rounds necessary to terminate. This is because while the small group size resulting from the arithmetic mapping gives rise to a small $M$, it also increases the number of samples required for elimination of $\beta$-suboptimal arms. Fig. 5(a) shows, the order between the cumulative regrets of the mappings as well as the communication costs. The geometric mapping offers the best performance with $M=92$ client, while the polynomial mapping offers the worst with the maximum number of clients $M=288$. Fig. 5(b) shows all mappings result in the termination of the algorithm with the geometric mapping giving the quickest termination and the arithmetic mapping giving the slowest.
@@ 6.3 Scenario 3
7 Conclusion and Future Work
@@
Appendix
Proof of Lemma 1
The proof builds upon the proof of MoM concentration bound given in [59, Proposition 2]. Let ${\mathcal{G}}_{p}$ denote the set of client groups formed by the server at the end of phase $p$ and ${\mathcal{G}}^{\text{tru}}_{p}\subseteq{\mathcal{G}}_{p}$ denote the subset of groups with no Byzantine clients. We denote by $G_{p}:=\lvert{\mathcal{G}}_{p}\rvert$ and $G^{\text{tru}}_{p}:=|{\mathcal{G}}^{\text{tru}}_{p}|$ the cardinalities of these sets. There is no concentration guarantee for the groups that include Byzantine clients. The worst case is when each Byzantine client belongs to a different group, rendering the maximum number of groups unreliable. Since $M^{\text{mal}}\leq\lambda M$ due to Assumption 1, at most $\lfloor\lambda M\rfloor$ of the $G_{p}$ groups will be corrupted by Byzantine clients. Then, we have,
$\mathbb{P}(|\overline{U}_{k}(p)-\mu_{k}|>\gamma_{p}|{\mathcal{G}}_{p})$
By definition, Term I is the average of $B$ independent $\sigma/\sqrt{s(p)}$-subgaussian random variables. Thus, Term I is $\sigma/\sqrt{s(p)B}$-subgaussian. Similarly, Term II is the average of $B$ independent $\sigma_{c}$-subgaussian random variables. Therefore, it is $\sigma_{c}/\sqrt{B}$-subgaussian. As a result, Term I plus Term II is $\sigma/\sqrt{s(p)B}+\sigma_{c}/\sqrt{B}$-subgaussian. Let $\rho\coloneqq\sigma/\sqrt{s(p)}+\sigma_{c}$ and $\rho^{\prime}\coloneqq\frac{1}{\sqrt{B}}\rho$. Then, using the subgaussian tail inequality in [63, Theorem 5.3 ]), we have
where (32) follows from (8), and (33) follows since $\eta(\lambda)>1/2$ implies that $4\eta(\lambda)-1>2\eta(\lambda)$. From (30), observe that (33) holds and $\eta(\lambda)G_{p}-G_{p}^{\text{tru}}\leq 0$, which gives
(30)$\leq q_{p}^{\frac{2\eta(\lambda)-1}{2}G_{p}}2^{\eta(\lambda)G_{p}}~{}.$ (34)
for all $\gamma_{p}^{2}\geq\frac{4\rho^{2}}{B}\frac{4\eta(\lambda)-1}{2\eta(\lambda)-1}\log 2$, which completes the proof.
Proof of Lemma 3
where (36) follows from the definitions of $\rho$ and $E_{p}$. Finally, observe that our choice of $M$ satisfies (36).